# Heights and periods of algebraic cycles in families

### v2

@@ 1.1. Main results on Gross–Schoen and Ceresa cycles
Let ${\mathscr{M}}_{g}$ be the moduli space of smooth projective curves of genus $g$, let ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$ be the universal curve, and let $\pi\colon\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g}){\longrightarrow}{\mathscr{M}}_{g}$ be the relative Picard group of degree $1$ on ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
@@
for $s=\pi(\xi)\in{\mathscr{M}}_{g}({\mathbb{C}})$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, a particularly important element $\xi_{s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ is such that the divisor class of $(2g-2)\xi_{s}$ is the canonical divisor class on ${\mathscr{C}}_{s}$; it is well-defined up to $(2g-2)$-torsions.
@@
For each $g\geq 3$, there exist a Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$, such that for any $s\in{\mathscr{M}}_{g}^{\mathrm{amp}}(\overline{\mathbb{Q}})$ and any $\xi\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$, we have
(1.1) $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{s}))-c,$
$\langle{{\mathrm{Ce}}(\xi),{\mathrm{Ce}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{s}))-c,$
@@
The proof will be executed in §$\S$6.3, in which we give a construction of ${\mathscr{M}}_{g}^{\mathrm{amp}}$. Indeed, we will take ${\mathscr{M}}_{g}^{\mathrm{amp}}$ to be ${\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, where ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is the Betti degeneracy locus (1.3) with the normal function $\nu$ being the Gross–Schoen normal function (C.6). A priori ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is only a real analytic set. But we show that ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is proper Zariski closed in ${\mathscr{M}}_{g,{\mathbb{C}}}$ by applying Ax-Schanuel multiple times; see Theorem 1.4 and Corollary 1.8. Finally, we show that it is defined over ${\mathbb{Q}}$ by volume calculation for adelic line bundles; see Corollary 6.4. Because of the importance of this Zariski open dense set ${\mathscr{M}}_{g}^{\mathrm{amp}}$ on which (1.1) holds, we will make the following definition: We call
@@
    ${\mathscr{M}}_{g}^{\mathrm{amp}}$ the ample locus of the Gross–Schoen and Ceresa cycles;
@@
    ${\mathscr{M}}_{g}^{\mathrm{slim}}:={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{amp}}$ the slim locus of the Gross–Schoen and Ceresa cycles. By the explanation above, it is the ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. It certainly contains the hyperelliptic locus and also the locus of Picard curves by Laga–Shnidman [LS24, Thm. C].
We expect ${\mathscr{M}}_{g}^{\mathrm{amp}}$ is the largest open subset of ${\mathscr{M}}_{g}$ so that the height inequality (1.1) in Theorem 1.1 holds; see Conjecture 1.12.
@@
In the course of the proof, we also establish the following geometric result on the torsion locus. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s}):={\mathrm{Ce}}(\xi_{s})$.
@@ 1.2.3. Consequence on torsion loci
The second part of Conjecture 1.12 is equivalent to: For any immersion $S\subseteq{\mathbb{P}}^{N}$ defined over $\overline{\mathbb{Q}}$, there exists a Zariski dense sequence $s_{i}\in S^{\mathrm{slim}}(\bar{K})$ such that
@@
The proof of Theorem 1.1 executed in §$\S$6.3 in fact proves the first part of Conjecture 1.12 in the case of Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$. A key ingredient is the volume identity (6.4). The same argument can be used to show that the following conjectured volume identity (1.10) implies the full Conjecture 1.12 with the help of [YZ21, Thm. 6.3.6]. To check whether $S^{\mathrm{amp}}$ is non-empty, one can use (1.6) and Theorem 1.6.
@@
(1.10) $\widehat{\mathrm{vol}}(\widetilde{L}|_{S^{\prime}})=\int_{S^{\prime}({\mathbb{C}})}c_{1}(\bar{L}_{v})^{\dim S}.$
@@ 1.4. Plan and Ingredients of proofs
The arithmetic part requires the construction of a suitable adelic line bundle $\overline{{\mathscr{L}}}$ over the base, a theory initiated by the second-named author [Zha95] and developed in further joint work with Yuan [YZ21], such that the height function $h_{\overline{{\mathscr{L}}}}$ is the Beilinson–Bloch height. Then, the desired height inequalities will follow from suitable bigness properties of $\overline{{\mathscr{L}}}$. In the context of Theorem 1.1 and Corollary 1.2, by the second-named author’s [Zha10, Thm. 2.5.5] the desired $\overline{{\mathscr{L}}}$ can be obtained from a suitable Deligne pairing. This $\overline{{\mathscr{L}}}$ has been constructed by Yuan [Yua21, §$\S$3.3.4] is slightly different steps. The required bigness property is the bigness of the generic fiber $\widetilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$. We furthermore prove the volume identity as Theorem 6.3, which implies that the bigness of $\widetilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$ with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form. The arithmetic part of this paper is confined to the Gross-Shoen and the Ceresa cycles.
@@ Acknowledgements
We would like to thank Ari Shnidman for asking us the question about the Northcott property for Ceresa cycles, Richard Hain for numerous inspiring discussions and for sharing his notes, Xinyi Yuan for his help during the preparation of this project, Dick Gross for suggesting we look at $\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ instead of just ${\mathscr{M}}_{g}$, and Peter Sarnak for asking the question about the positivity of the Beilinson–Bloch heights. Part of this work was completed when ZG visited SZ at Princeton University in January, February, and May 2024. ZG would like to thank Princeton University for its hospitality. We also appreciate your comments on an earlier draft.
@@ 3.1. Replacing $S$ by $\varphi(S)$
Recall from Remark 2.2 that ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$ is an intermediate Jacobian over $S$. Set ${\mathscr{J}}^{\prime}:=S^{\prime}\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$; it is an intermediate Jacobian over $S^{\prime}$. Then the inclusion $S^{\prime}\subseteq\Gamma\backslash{\mathscr{D}}$ yields a section $\nu^{\prime}$ of ${\mathscr{J}}^{\prime}\rightarrow S^{\prime}$, and thus we can define $S^{\prime}_{{\mathscr{F}}}(t)$ for each $t\geq 0$ with respect to $\nu^{\prime}$. We have the following commutative diagram, with $\varphi_{{\mathscr{J}}}$ induced by $\varphi$, such that $\varphi_{{\mathscr{J}}}\circ\nu=\nu^{\prime}\circ\varphi$:
@@
(3.3) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime}_{{\mathscr{F}}}(t-r)\right).$
@@ 3.3. Applications of Ax–Schanuel
Let us prove $\geq$”. Let $t=\dim S-r(\nu)$. Then $S^{\mathrm{Betti}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
@@ 3.5. Zariski closedness of the degeneracy loci
By (3.3), Proposition 3.9 and Lemma 3.10 (both applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
@@ 6.1. Construction of the adelic line bundle
The following theorem is a reformulation of the second-named author’s [Zha10, Thm. 2.3.5]. Recall the notation ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{s})$ with $\xi_{s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ such that $(2g-2)\xi_{s}$ being the canonical divisor class on ${\mathscr{C}}_{s}$.
@@ 6.3. Bigness of the generic fiber of the adelic line bundle
Let $S^{\prime}$ be an irreducible component of $S^{\mathrm{Betti}}(1)$. Then $S^{\prime}=S^{\prime}_{{\mathscr{F}}}(1)$, and hence $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime}$; see (1.6) and Corollary D.7. So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S^{\prime}})=0$ by Theorem 6.3. By Definition E.11, for any automorphism $\sigma\in{\mathrm{Aut}}({\mathbb{C}}/k)$ we have
@@
$S^{\prime\sigma}({\mathbb{C}})=S^{\prime\sigma}_{\mathscr{F}}(1)\subseteq S^{\mathrm{Betti}}(1).$
@@
Proof of Theorem 1.1.
It suffices to prove the claim for $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}$ because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple.
Next by [Zha10, Cor. 2.5.2], it suffices to prove for $\xi=\xi_{s}$. We start by showing there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$ such that
@@
for all $s\in U(\overline{\mathbb{Q}})$. The conclusion follows immediately from Theorem 6.5 and [YZ21, Thm. 5.3.5.(iii)] (which we reformulated as Theorem E.9). Notice that one can either take the adelic line bundle on ${\mathscr{M}}_{g}$, which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
We can do better to get an explicit $U$. Suppose $U_{1},U_{2}$ are open subsets of ${\mathscr{M}}_{g}$ over ${\mathbb{Q}}$ such that (6.5) holds true for all $s\in U_{1}(\overline{\mathbb{Q}})$ and $s\in U_{2}(\overline{\mathbb{Q}})$, then (6.5) holds true for all $s\in(U_{1}\cup U_{2})(\overline{\mathbb{Q}})$. Therefore, there exists a maximal Zariski open dense subset $U^{\prime}$ of ${\mathscr{M}}_{g}$ on which (6.5) holds. Let $S:={\mathscr{M}}_{g}\setminus U^{\prime}$. Then $S$ is proper Zariski closed in ${\mathscr{M}}_{g}$, and the height inequality (6.5) cannot hold for any non-empty open subvariety $V$ of $S$. By Theorem E.9, $\widetilde{{\mathscr{L}}}$ cannot be big on any irreducible component $Z$ of $S$, and hence $c_{1}(\overline{\mathscr{L}}|_{Z({\mathbb{C}})})^{\wedge\dim Z}\equiv 0$ by Proposition 6.3. Thus $Z_{{\mathbb{C}}}=Z_{{\mathscr{F}}}(1)\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. So $S_{{\mathbb{C}}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. So $S\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ by Corollary 6.4. Thus we can take ${\mathscr{M}}_{g}^{\mathrm{amp}}$ to be ${\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, which is Zariski open dense and is defined over ${\mathbb{Q}}$, in Theorem 1.1.∎

### v3

@@ 1.1. Main results on Gross–Schoen and Ceresa cycles
Let ${\mathscr{M}}_{g}$ be the coarse moduli space of smooth projective curves of genus $g$. Let ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$ be the universal curve, and let $\pi\colon\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g}){\longrightarrow}{\mathscr{M}}_{g}$ be the relative Picard group of degree $1$ on ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$, both as Deligne–Mumford stacks. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
@@
for $s=\pi(\xi)\in{\mathscr{M}}_{g}({\mathbb{C}})$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, a particularly important element $\xi_{\mathrm{can},s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ is such that the divisor class of $(2g-2)\xi_{\mathrm{can},s}$ is the canonical divisor class on ${\mathscr{C}}_{s}$; it is well-defined up to $(2g-2)$-torsions.
@@
For each $g\geq 3$, there exist a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$, such that for any $s\in U(\overline{\mathbb{Q}})$ and any $\xi\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$, we have
(1.1) $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{\mathrm{can},s}))-c,$
$\langle{{\mathrm{Ce}}(\xi),{\mathrm{Ce}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{\mathrm{can},s}))-c,$
@@
If (1.1) holds over Zariski open dense subsets $U_{1}$ and $U_{2}$ of ${\mathscr{M}}_{g}$, then it holds also over $U_{1}\cup U_{2}$. Thus, we can propose the following definition:
@@
    the ample locus of the Gross–Schoen and Ceresa cycles, denoted by ${\mathscr{M}}_{g}^{\mathrm{amp}}$, is the maximal open subset of ${\mathscr{M}}_{g}$ over ${\mathbb{Q}}$ so that (1.1) holds;
@@
    ${\mathscr{M}}_{g}^{\mathrm{slim}}:={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{amp}}$ the slim locus of the Gross–Schoen and Ceresa cycles.
Theorem 1.1 is then equivalent to: ${\mathscr{M}}_{g}^{\mathrm{amp}}$ is non-empty if $g\geq 3$. The proof will be executed in §$\S$6.3, in which we give an explicit construction of ${\mathscr{M}}_{g}^{\mathrm{amp}}$ (or equivalently its complement ${\mathscr{M}}_{g}^{\mathrm{slim}}$). Indeed, we will show that ${\mathscr{M}}_{g}^{\mathrm{slim}}$ is precisely the Betti degeneracy locus ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ defined in (1.3) with the normal function $\nu$ being the Gross–Schoen normal function (C.6). A priori ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is only a real analytic set. But we show that ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is proper Zariski closed in ${\mathscr{M}}_{g,{\mathbb{C}}}$ by applying Ax-Schanuel multiple times; see Theorem 1.4 and Corollary 1.8. Finally, we show that it is defined over ${\mathbb{Q}}$ by volume calculation for adelic line bundles; see Corollary 6.4. We point out that ${\mathscr{M}}_{g}^{\mathrm{slim}}$ contains the hyperelliptic locus and also the locus of Picard curves by Laga–Shnidman [LS24, Thm. C].
@@
In the course of the proof, we also establish the following geometric result on the torsion locus. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{\mathrm{can},s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s}):={\mathrm{Ce}}(\xi_{\mathrm{can},s})$.
@@ 1.2.3. Consequence on torsion loci
In practice, in addition to this conjecture one also needs to check whether $S^{\mathrm{amp}}$ is empty, by using (1.6) and Theorem 1.6. The second part of Conjecture 1.12 is equivalent to: For any immersion $S\subseteq{\mathbb{P}}^{N}$ defined over $\overline{\mathbb{Q}}$, there exists a Zariski dense sequence $s_{i}\in S^{\mathrm{slim}}(\bar{K})$ such that
@@
The proof of Theorem 1.1 executed in §$\S$6.3 in fact proves the following:
Theorem 1.13.
Conjecture 1.12 holds in the case of Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$.
A key ingredient to prove this theorem is the volume identity (6.4). The same argument can be used to show that the following conjectured volume identity (1.10) implies the full Conjecture 1.12 with the help of [YZ21, Thm. 5.3.8].
@@
(1.10) $\widehat{\mathrm{vol}}(\widetilde{L}|_{S^{\prime}})=\int_{S^{\prime}({\mathbb{C}})}c_{1}(\bar{L}_{v})^{\wedge\dim S^{\prime}}.$
@@ 1.4. Plan and Ingredients of proofs
The arithmetic part requires the construction of a suitable adelic line bundle $\overline{{\mathscr{L}}}$ over the base, a theory initiated by the second-named author [Zha95] and developed in further joint work with Yuan [YZ21], such that the height function $h_{\overline{{\mathscr{L}}}}$ is the Beilinson–Bloch height. Then, the desired height inequalities will follow from suitable bigness properties of $\overline{{\mathscr{L}}}$. In the context of Theorem 1.1 and Corollary 1.2, by the second-named author’s [Zha10, Thm. 2.5.5] the desired $\overline{{\mathscr{L}}}$ can be obtained from a suitable Deligne pairing. This $\overline{{\mathscr{L}}}$ has been constructed by Yuan [Yua21, §$\S$3.3.4] in slightly different steps, where he also asked the question of nefness and bigness of $\overline{{\mathscr{L}}}$. In our paper, the required bigness property is the bigness of the generic fiber $\widetilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$. We furthermore prove the volume identity as Theorem 6.3, which implies that the bigness of $\widetilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$ with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form. The arithmetic part of this paper is confined to the Gross-Shoen and the Ceresa cycles.
@@ Acknowledgements
We would like to thank Ari Shnidman for asking us the question about the Northcott property for Ceresa cycles, Richard Hain for numerous inspiring discussions and for sharing his notes, Xinyi Yuan for his help during the preparation of this project and comments on a previous version of the paper, Dick Gross for suggesting we look at $\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ instead of just ${\mathscr{M}}_{g}$, and Peter Sarnak for asking the question about the positivity of the Beilinson–Bloch heights. Part of this work was completed when ZG visited SZ at Princeton University in January, February, and May 2024. ZG would like to thank Princeton University for its hospitality. We also appreciate your comments on an earlier draft.
@@ 3.1. Replacing $S$ by $\varphi(S)$
Recall from Remark 2.2 that ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$ is an intermediate Jacobian over $S$. Set ${\mathscr{J}}^{\prime}:=S^{\prime}\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$; it is an intermediate Jacobian over $S^{\prime}$. Then the inclusion $S^{\prime}\subseteq\Gamma\backslash{\mathscr{D}}$ yields a section $\nu^{\prime}$ of ${\mathscr{J}}^{\prime}\rightarrow S^{\prime}$, and thus we can define $S^{\prime\mathrm{Betti}}(t)$ for each $t\geq 0$ with respect to $\nu^{\prime}$. We have the following commutative diagram, with $\varphi_{{\mathscr{J}}}$ induced by $\varphi$, such that $\varphi_{{\mathscr{J}}}\circ\nu=\nu^{\prime}\circ\varphi$:
@@
(3.3) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime\mathrm{Betti}}(t-r)\right).$
@@ 3.3. Applications of Ax–Schanuel
Let us prove $\geq$”. Let $t=\dim S-r(\nu)$. Then $S^{\mathrm{Betti}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime\mathrm{Betti}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
@@ 3.5. Zariski closedness of the degeneracy loci
By (3.3), Proposition 3.9 and Lemma 3.10 (both applied to $S^{\prime\mathrm{Betti}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
@@ 6.1. Construction of the adelic line bundle
The following theorem is a reformulation of the second-named author’s [Zha10, Thm. 2.3.5]. Recall the notation ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{\mathrm{can},s})$ with $\xi_{\mathrm{can},s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ such that $(2g-2)\xi_{\mathrm{can},s}$ being the canonical divisor class on ${\mathscr{C}}_{s}$.
@@ 6.3. Bigness of the generic fiber of the adelic line bundle
Let $S^{\prime}$ be an irreducible component of $S^{\mathrm{Betti}}(1)$. Then $S^{\prime}=S^{\prime\mathrm{Betti}}(1)$, and hence $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime}$; see (1.6) and Corollary D.7. So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S^{\prime}})=0$ by Theorem 6.3. By Definition E.11, for any automorphism $\sigma\in{\mathrm{Aut}}({\mathbb{C}}/k)$ we have
@@
$S^{\prime\sigma}({\mathbb{C}})=(S^{\prime\sigma})^{\mathrm{Betti}}(1)\subseteq S^{\mathrm{Betti}}(1).$
@@
Proof of Theorem 1.1 and of ${\mathscr{M}}_{g}^{\mathrm{amp}}={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
We start by proving Theorem 1.1. It suffices to prove the height comparison (1.1) for $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}$ because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple. Next by [Zha10, Cor. 2.5.2], it suffices to prove (1.1) for $\xi=\xi_{\mathrm{can},s}$, i.e. there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$ such that
@@
for all $s\in U(\overline{\mathbb{Q}})$. The conclusion follows immediately from Theorem 6.5 and [YZ21, Thm. 5.3.5.(iii)] (which we reformulated as Theorem E.9). Notice that one can either take the adelic line bundle $\overline{{\mathscr{M}}}$ on ${\mathscr{M}}_{g}$ which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
Now we prove ${\mathscr{M}}_{g}^{\mathrm{amp}}={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, or equivalently ${\mathscr{M}}_{g}^{\mathrm{slim}}={\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. Notice that ${\mathscr{M}}_{g}^{\mathrm{slim}}$ is by definition defined over ${\mathbb{Q}}$ and ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is defined over ${\mathbb{Q}}$ by Corollary 6.4.
We start by showing ${\mathscr{M}}_{g}^{\mathrm{slim}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. By definition of ${\mathscr{M}}_{g}^{\mathrm{slim}}$, the height inequality (6.5) cannot hold for any non-empty open subvariety $V$ of ${\mathscr{M}}_{g}^{\mathrm{slim}}$. By Theorem E.9, $\widetilde{{\mathscr{L}}}$ cannot be big on any irreducible component $S$ of ${\mathscr{M}}_{g,\overline{\mathbb{Q}}}^{\mathrm{slim}}$, and hence $c_{1}(\overline{\mathscr{L}}|_{S({\mathbb{C}})})^{\wedge\dim S}\equiv 0$ by Proposition 6.3. Thus $S=S^{\mathrm{Betti}}(1)\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. So ${\mathscr{M}}_{g,\overline{\mathbb{Q}}}^{\mathrm{slim}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. So ${\mathscr{M}}_{g}^{\mathrm{slim}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ by Corollary 6.4.
It remains to prove that $S:={\mathscr{M}}_{g}^{\mathrm{Betti}}(1)\setminus{\mathscr{M}}_{g}^{\mathrm{slim}}={\mathscr{M}}_{g}^{\mathrm{Betti}}(1)\cap{\mathscr{M}}_{g}^{\mathrm{amp}}$ is empty. By definition of ${\mathscr{M}}_{g}^{\mathrm{amp}}$, the inequality (6.5) holds for all $s\in S(\overline{\mathbb{Q}})$ and $S$ is Zariski open in ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. If $S\not=\emptyset$, then ~$\widetilde{{\mathscr{L}}}|_{Z}$ is big by [YZ21, Thm. 5.3.8] for some irreducible component $Z$ of ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, and hence $c_{1}(\overline{{\mathscr{L}}}|_{Z({\mathbb{C}})})^{\dim Z}\not\equiv 0$. But this contradicts the construction of ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$. So $S=\emptyset$, and we are done. ∎